A new movie premiered on Friday September 2, and 1350 people attended.Attendance then decreased by 20% each day. Write an exponential decay function to model this situation.Then find the attendance on September 7
Answer by JoelSchwartz(130)

y=1350
y-.2y=1,080
1,080-.2*1,080=864
z=percent of 1350 that equals 864
1350z=864
z=864/1350
z=.64
.8y=1,080
.64y=864
x=number of days after September second
1350*.8^x
1350*.8^(5)=442.368
The attendance on September seven would be four hundred forty two people.